IODP Proceedings    Volume contents     Search



The D/V Chikyu moved toward Site C0002 after midnight on 12 November 2010, moving upstream of the site to prepare the 20 inch casing and jetting-in the bottom-hole assembly (BHA) with the mudmat and wellhead assembly connected (Table T1). At 1830 h on 13 November, the BHA was run into the water ~15 nmi upstream of Site C0002. The remotely operated vehicle (ROV) dove at about the same time, and the BHA was run into the water, with 40 m of the 20 inch casing hanging below, until it reached 1441 m drilling depth below rig floor (DRF), whereupon anti-vortex-induced vibration (VIV) ropes were attached to the upper portion of the drill pipe as it continued being run. By 0330 h on 14 November, the BHA reached 1861 m DRF while the Chikyu continued to drift to Site C0002. The Chikyu reached 2.5 nmi from Site C0002 by 0600 h on 15 November while the ROV continued monitoring the BHA and running tool for signs of VIV. At 1247 h the BHA position was confirmed by ROV and the ship’s dynamic positioning (DP) systems, and the seafloor was tagged at 1256 h. Hole C0002G was at 33°18.0126′N, 136°1488′E in water depth of 1966 m DRF (1937.5 m mud depth below sea level). Jetting in with the 20 inch casing began at 1307 h. The cart was released by the ROV at 1630 h, and beginning at 1700 h the running tool was picked up to 1970 m DRF while ropes were removed from the drill string.

The BHA was pulled out of the hole to the surface and was laid down at 0330 h on 16 November. Once recovered, there was a short stop at 0710 h as the Chikyu moved to a low-current area (LCA) to rendezvous with the Heisei maru to backload the circulation obviation retrofit kit (CORK) running tool for transport ashore. At 1230 h the guide horn was moved to the moonpool for assembly and was completed by 1815 h, after which the 12¼ inch drilling LWD/MWD BHA was made up and run into the hole at 1900 h. By 0618 h on 17 November, the BHA had reached 1787 m DRF as the Chikyu drifted along with the current, reentering the wellhead at 0945 h. Reaming to the bottom of the 20 inch casing at 2007 m DRF began, and drilling with LWD to 2267 m DRF commenced at 1030 h. Drilling reached 2573 m DRF at 1145 h on 18 November when a short trip to the casing shoe to circulate and clean the hole was run from 1230 to 1430 h. The run back to the bottom resumed at 1630 h, and drilling resumed from 2573 to 2675 m DRF at 1830 h. The BHA reached 2858 m DRF, an initial 9⅝ inch casing shoe target depth, at 1045 h on 19 November when cleaning and sweeping of the hole was resumed.

The BHA was pulled out of the hole to 1810 m DRF with no overpull and then resumed pulling out of the hole to the rig floor. The ROV was recovered at 1640 h as the Chikyu drifted along the current and continued to pull the BHA out of the hole. The LWD/MWD BHA was laid down at 0100 h on 20 November, and 10⅝ inch BHA was made up and run into the hole at 0130 h. At 0800 h, 1 nmi from Hole C0002G, the ROV dove and the BHA run into the hole was paused, resuming at 1030 h, and the BHA reentered the hole at 1130 h. By 1145 h the BHA was at 2785 m DRF, where a short relogging trip began from 2785 to 2804 m DRF and again from 2830 to 2858 m DRF, first to overlap the proposed casing shoe depth and also to confirm the suitability of the proposed strainmeter depth and sensor carrier positioning. At 1900 h, drilling with the 10⅝ inch BHA LWD assembly began from 2858 m DRF to total depth (TD) (2946 m DRF), which was reached at 0230 h on 21 November when circulation to clean the hole began, sweeping with 10 m3 of Hi-Vis mud. After sweeping was completed at 0445 h, the BHA was pulled out of the hole to 2862 m DRF to relog from 2859 to 2830 m DRF because the live-streaming LWD data were missing two resistivity channels and did not appear in play back after TD was reached. Relogging finished at 0600 h, after which the BHA made a short trip to 2500 m DRF.

The BHA was run down to the bottom at 0715 h for circulation. There was some concern about hole conditions deteriorating below 2934 m DRF, as the BHA began experiencing pack off. Circulation began at 0900 h, and a wiper trip to the 20 inch casing shoe was completed at 1200 h. A run back to the bottom, starting at 1500 h, showed no more trouble with hole conditions, as reaming slowly to 2945 m DRF had no problems. Hi-Vis mud was spotted at 1900 h, pulling out of the hole to 1870 m DRF began at 2015 h, and the BHA was clear of the wellhead at 2330 h. By 0615 h on 22 November, the 10⅝ inch BHA was laid down, and then the 9⅝ inch casing was readied to be run in the hole while moving to the LCA. At 0845 h the guide horn was rigged down, and at 1130 h the casing was rigged up and ready to run. At 1330 h the casing was run in, reaching 130 m DRF by 1730 h. The casing screen joints were run in the hole from 1330 to 1945 h, followed by the rest of the casing, while the Chikyu began drifting back to Hole C0002G.

Anti-VIV ropes were deployed on the drill pipe from 1360 m DRF at 0600 h on 23 November until 1830 m DRF was reached at 2000 h that evening. Once the ropes were securely tied off (terminated) on the drill string, the guide horn was reinstalled and completed by 2300 h. From that point, drifting back to reenter Hole C0002G continued all day on 24 November, reaching 1.5 nmi from the location at 2300 h, whereupon the ROV dove. No VIV was observed throughout the entire time the drill string was hung while drifting back to Hole C0002G.

The casing BHA reached 10 m above the wellhead at 0300 h on 25 November, and the ROV stood by to observe reentry into the wellhead. The casing was run into the wellhead, reaching the planned setting depth (2853 m DRF) by 0530 h. Cementing the casing began at 0930 h on 25 November, after which the casing hanger tool was released and pulled out of the hole while the drill string was cleared. The ROV was recovered on deck by 1330 h, and the casing hanger was recovered on deck by 0200 h on 26 November. The cement stand and head were examined, and the wiper plug was found trapped above the drop indicator inside the cement head. The Chikyu began moving to Shimizu, Japan, for port call, arriving by 0700 h on 27 November.

The Chikyu stayed quayside, loading and unloading tools and equipment, leaving port at 2000 h and making way back to Site C0002. The Chikyu reached 7 nmi upstream of Site C0002 at 1730 h on 28 November, went into DP mode, and installed the ROV cursor rail at 2030 h. The drill out cement BHA was run into the hole to 1860 m DRF by 0130 h on 29 November while drifting back toward Hole C0002G.

The ROV dove at 0315 h on 29 November, and at 0830 h the BHA was run into the hole to 2826 m DRF to begin drilling out cement. After a short trip to the 9⅝ inch float shoe, the BHA was run to 2943 m DRF, the hole was swept, and high-density brine and protect zone were spotted into the hole at 1530 h. After spotting, the BHA was pulled out of the hole from the wellhead at 1715 h, and then the Chikyu drifted as the drill pipe continued to be recovered.

Taking position 20 nmi northwest of Site C0002, the CORK running tool was picked up and makeup began at 0500 h on 30 November. The upper guide horn was rigged down at 0730 h, and the completion assembly was prepared for running into the hole. Running into the hole began at 2300 h on 30 November, with stopping for connection and termination of three cables, the flatpack, and the thermistor string on the sensor carrier instruments at 0700 h on 1 December. Joints of 3½ inch tubing were run with four centralizers per joint up to the swellable packer at 1400 h on 2 December. The electrical cables, flatpack, and hydraulic lines were run through the packer body seal by 1630 h on 2 December, after which the electrical cables were tested at 2215 h to ensure that the connections to the sensors were still viable after working them through the packer.

Running continued from 2300 h on 2 December to 948 m DRF on 3 December, when the CORK head was readied for attachment to the completion assembly. The CORK running tool was picked up and mated to the CORK head by 0045 h on 4 December and run through the rotary table by 0200 h. The electrical cables were again tested at 0400 h, after which the cables were measured for cutting and termination. Termination by Ocean Design, Inc. (ODI) engineers began at 0730 h on 4 December, and the terminated cables were installed on the CORK head at 0830 h on 5 December. The ROV platform was lowered to the moonpool in preparation for assembly and attachment to the CORK head at 1645 h. Stands of drill pipe were fixed to the CORK running tool and run into the hole to 1300 m DRF by 2230 h on 5 December while the ROV entered the water. The completion assembly was run to 1417 m DRF, where the ROV tested the underwater mateable connectors (UMCs) on the CORK head at 0200 h on 6 December. All connections passed inspection, and the completion assembly continued being run into the hole to 1897 m DRF by 0045 h on 8 December. The UMCs were successfully checked again by ROV at 0430 h on 8 December before running the CORK assembly into the wellhead.

The Chikyu approached and shifted position before running the CORK completion assembly into the wellhead at 0845 h. The assembly was run to 2904 m DRF, and the CORK head was landed on the wellhead at 2015 h on 8 December. One more systems check was performed on the tiltmeter and strainmeter UMCs only because the broadband seismometer connector was in danger of being damaged by the ROV arm. All checks were successfully completed by 2200 h. Cementing was finished by 0230 h on 9 December, and the unjay stopper and CORK running tool were released at 0330 h. The running tool was pulled out of the hole to the surface by 1355 h on 9 December.